# NovelSeek: When Agent Becomes the Scientist – Building Closed-Loop System from Hypothesis to Verification

### v2

@@ Abstract
Artificial Intelligence (AI) is accelerating the transformation of scientific research paradigms, not only enhancing research efficiency but also driving innovation. We introduce NovelSeek, a unified closed-loop multi-agent framework to conduct Autonomous Scientific Research (ASR) across various scientific research fields, enabling researchers to tackle complicated problems in these fields with unprecedented speed and precision. NovelSeek highlights three key advantages: 1) Scalability: NovelSeek has demonstrated its versatility across 12 scientific research tasks, capable of generating innovative ideas to enhance the performance of baseline code. 2) Interactivity: NovelSeek provides an interface for human expert feedback and multi-agent interaction in automated end-to-end processes, allowing for the seamless integration of domain expert knowledge. 3) Efficiency: NovelSeek has achieved promising performance gains in several scientific fields with significantly less time cost compared to human efforts. For instance, in reaction yield prediction, it increased from 27.6% to 35.4% in just 12 hours; in enhancer activity prediction, the Pearson correlation coefficient rose from 0.65 to 0.79 with only 4 hours of processing; and in 2D semantic segmentation, precision advanced from 78.8% to 81.0% in a mere 30 hours.
Figure 1: NovelSeek can support 12 types of scientific research tasks ranging from the AI field to the science field, including reaction yield prediction, molecular dynamics, power flow estimation, time series forecasting, transcription prediction, enhancer activity prediction, sentiment classification, 2D image classification, 3D point classification, 2D semantic segmentation, 3D autonomous driving, large vision-language model fine-tuning.
@@ 1 Introduction
To further facilitate the advancement of ASR, we propose the NovelSeek, an end-to-end auto-research pipeline, which covers four main modules: self-evolving idea generation, human-interactive feedback, idea-to-methodology construction, and multi-round experiment planning and execution. With the help of the Self-Evolving Human-interactive Idea Generation and Idea-to-Methodology Construction, NovelSeek can transform a rough proposal into a detailed and easily implementable method, which further increases the success rate of code implementation process and enhances the efficiency of closed-loop experiments. Besides, by leveraging multi-round experiment planning and execution, NovelSeek is capable of designing experimental plans and decomposing the experimental process according to the NovelSeek-proposed modules, thereby validating the effectiveness of each NovelSeek-generated module through experimentation.
As shown in Fig. 1, NovelSeek has been validated across 12 scientific research tasks, and we are excited to see that the experimental results demonstrate the significant value of NovelSeek in the entire process from hypothesis generation to experimental validation. For instance, in the Reaction Yield Prediction task, the baseline model only achieved a performance of $24.2\%\pm 4.2$, while our model improved it to $34.8\%\pm 1.1$ in just 12 hours. In contrast, human researchers typically require several months to achieve a similar level of performance improvement. Another example of performance improvement is the Enhancer Activity Prediction task. The baseline model, DeepSTARR, achieved a result of 0.65. By utilizing NovelSeek to search relevant domain literature, automatically generate code, and conduct validation, the performance can be improved to 0.79, representing a promising enhancement. In addition, NovelSeek also supports complex project-level modifications and debugging, which consist of multiple code files. These results clearly indicate that NovelSeek can autonomously generate ideas and design algorithms, effectively reducing the dependence on human effort in scientific research. To facilitate reproducibility, we have open-sourced both the baselines and the codes generated by NovelSeek used in all involved scientific research tasks at https://github.com/Alpha-Innovator/NovelSeek.
@@
  • Unified Multi-agent Framework for Diverse Scientific Research Tasks: We present NovelSeek, a unified closed-loop scientific research framework that can automate the entire research cycle, including idea generation, idea-to-methodology transformation, experiment execution, and result feedback. This unified framework can be directly applied to various scientific research scenarios and fields.
  • Interactive Interfaces for Enhanced Cooperative Research: NovelSeek offers interactive interfaces for human-machine collaboration within the idea generation module and across the entire system. By selecting collaboration modes, such as leveraging AI or human experts, it provides evaluations of idea generation effectiveness and facilitates the assessment, reflection, and documentation of experimental results.
  • Comprehensive Experimental Validation and Human Studies: We conducted extensive human studies centered around NovelSeek, including inviting domain experts to evaluate and score the novelty of ideas generated by NovelSeek, and comparing the research efficiency between human researchers and NovelSeek. These experiments and human studies are crucial for gaining insights into the capabilities of multi-agent systems in conducting scientific research tasks in open-ended environments. We observed many promising phenomena, while also identifying certain technical modules that require improvement.
Figure 2: NovelSeek covers three main capabilities: 1) Self-evolving Idea Generation with Human-interactive Feedback, 2) Idea-to-Methodology Construction, and 3) Evolutionary Experimental Planning and Execution.
2 NovelSeek
As a unified closed-loop multi-agent framework for ASR, NovelSeek is designed to facilitate innovative research across diverse scientific domains, as illustrated in Fig. 2. It incorporates three primary capabilities: self-evolving idea generation with human-interactive feedback (Sec. 2.1), comprehensive idea-to-methodology construction (Sec. 2.2), and multi-round automated experiment execution (Sec. 2.3). Each capability is realized through the collaboration of specialized agents, allowing for seamless integration of different processes to enhance scientific discoveries.
@@ 2.1 Self-Evolving Idea Generation with Human-interactive Feedback
The self-evolving idea generation capability is central to NovelSeek, enabling the framework to autonomously generate and refine innovative research ideas. This process involves several specialized agents, each contributing to different stages of idea development and refinement.
@@
By dynamically adjusting its search strategies based on the context of the research stage, the Survey Agent ensures a comprehensive and nuanced understanding of the research landscape. This capability not only supports the generation of innovative ideas but also ensures that the NovelSeek framework remains at the cutting edge of scientific discovery.
Code Review Agent. The Code Review Agent is crucial for understanding baseline codes for different research tasks, serving as a foundation for innovation by identifying improvements and developing new methodologies. It provides detailed analyses of code structures, dependencies, and functionalities, enabling NovelSeek to fully comprehend existing code-bases and identify potential enhancements to advance research objectives. Moreover, the agent’s ability to document and summarize complex code-bases ensures efficient navigation and utilization of existing methods. The agent manages two scenarios: 1) reviewing user-provided code or 2) searching for relevant code-bases. For user-uploaded code, it conducts a comprehensive review of the structure, logic, and functionality. Alternatively, in the absence of user-uploaded code, it searches public repositories like GitHub to find relevant code-bases, performing thorough analyses at both the repository and file levels to understand inter-dependencies and assess logic, efficiency, and correctness. Furthermore, the agent employs static code analysis using Python’s ‘ast‘ module to parse and understand code structure without execution, while the LLM generates human-readable descriptions and summaries, transforming technical details into structured documentation. By using parallel processing with Python’s ‘multiprocessing‘ module, the agent enhances efficiency and scalability for large code-bases. Overall, the Code Review Agent offers detailed documentation that deepens the understanding of code repositories and supports innovation in scientific research.
Idea Innovation Agent. The Idea Innovation Agent is an integral part of NovelSeek, designed to enhance the creative and iterative processes of scientific research. This agent plays a crucial role by automating the generation and evolution of ideas, thereby addressing the limitations of traditional research works (Yuan et al., 2025; Yamada et al., 2025; Lu et al., 2024), which often rely on time-consuming manual efforts and are constrained by human cognitive biases. The agent’s dual responsibilities, idea generation and idea evolution, are specifically designed to address the diverse needs of various scientific disciplines.
@@
Overall, the Idea Innovation Agent enhances scientific ideas into viable and creative solutions by synthesizing and contextualizing information. It critically examines current ideas and employs feedback loops with human experts and other NovelSeek agents for continuous improvement. This iterative process balances novelty, feasibility, and ethical considerations, producing impactful and well-rounded ideas.
Assessment Agent. The Assessment Agent is a vital component of NovelSeek, designed to ensure the quality and viability of generated ideas through a rigorous evaluation process. In the rapidly evolving landscape of scientific research, the systematic and objective assessment of ideas is essential. Traditional methods often suffer from subjectivity and lack comprehensive coverage of all relevant dimensions, which can lead to promising ideas being overlooked(Qiu et al., 2025; Si et al., 2024). Therefore, the Assessment Agent addresses these challenges by providing a structured and multidimensional evaluation process, which in turn enhances the reliability and effectiveness of idea selection.
@@
The human-interactive feedback mechanism of NovelSeek is categorized into two primary types: 1) feedback directly provided by humans and 2) feedback automatically generated by agent. Human-provided feedback can address one or multiple ideas, offering insights and critiques that lead to further refinement and adjustment of these ideas based on the feedback received. This iterative process facilitates the continuous improvement of ideas, ensuring they are honed to meet specific objectives and challenges.
@@
In summary, as illustrated in Fig. 3, by managing multi-agent collaboration and integrating human feedback, the Orchestration Agent enables NovelSeek to operate as a cohesive and innovative research tool, driving scientific discovery forward.
Figure 3: NovelSeek Self-evolutionary path of ideas for reaction yield prediction task.
@@ 2.2.2 Methodology Refinement
During both initialization and refinement, the Methodology Development Agent collaborates closely with other agents, such as the Assessment Agent for multidimensional evaluation and the Orchestration Agent for workflow coordination. This collaboration ensures that each methodological step benefits from comprehensive feedback and current domain knowledge. The integrated, multi-agent approach guarantees that the transformation from idea to methodology is systematic and adaptable, supporting the continuous evolution and optimization of scientific research within the NovelSeek framework.
@@ 3 Experiments
In this section, we evaluate the effectiveness of NovelSeek in conducting autonomous research and accelerating scientific discovery. We begin by providing a brief overview of the selected multi-domain tasks and detailing the experimental implementation in Sec. 3.1. Subsequently, we present the quantitative results across various tasks in Sec. 3.2 and conduct an analysis of the different modules within NovelSeek in Sec. 3.3.
@@ 3.1.1 Task Description
We select 12 distinct tasks to demonstrate NovelSeek’s capability in conduct Autonomous Scientific Research (ASR). These tasks span multiple modalities, including science (e.g., reaction yield prediction, molecular dynamics), time series (e.g., time series forecasting), natural language (e.g., sentiment classification), image (e.g., semantic segmentation), and point cloud (e.g., 3D object detection), which cover both discriminative and generative tasks. We believe that experiments ranging from fundamental tasks to complex multi-modal tasks can comprehensively illustrate the effectiveness of NovelSeek. Below, we detail the datasets, the base code repositories, and the experimental settings for each task.
@@ 3.1.2 Evaluation Metric
Since our NovelSeek has been validated across a wide range of scientific research fields, the evaluation metrics used for tasks in each field are not consistent. Therefore, in this part, we provide a detailed introduction to the evaluation metrics used for each scientific research task.
@@ 3.2 Experimental Results
Table 1: Performance comparison across six types of scientific research tasks. We conduct experiments using 10 NoveSeek generated ideas for each task.
@@
NovelSeek 33.5 (+5.9) 0.152 0.00447 0.4346 0.170 0.77
Table 2: Performance comparison for six types of scientific research tasks. We conduct experiments using 10 NoveSeek generated ideas for each task, where baseline codes for Auto2DSeg, AutoPCDet, and AutoVLM are project-level, consisting of multiple code files with complex call relation between functions. Therefore, the coder in Dolphin (Yuan et al., 2025) does not support modifying this type of baseline codes.
@@
NovelSeek 5/7/10 3/6/10 9/9/10 6/9/10 2/5/10 1/5/10
To comprehensively evaluate the effectiveness of NovelSeek in accelerating scientific discovery, we first provide quantitative experimental results as shown in Tab. 1, Tab. 2, Tab. 3, and Tab. 4. Extensive results demonstrate that NovelSeek excels in the following aspects:
Table 5: Computational and financial cost analysis for all tasks. Training time is measured using A100 GPU hours, while idea generation and code debugging costs are measured in USD using gpt-4o and claude-sonnet-3.7 models respectively.
@@
Coder-Debug cost (claude-sonnet-3.7) (\$) 0.7 0.5 1.0 0.4 0.9 0.6
@@
  • Outperforming existing auto-research systems on multiple tasks. We first compare NovelSeek with existing auto-research system (i.e., Dolphin (Yuan et al., 2025)) on single-file tasks. Tab. 1 and Tab. 2 show the max performance and average performance (i.e., the average performance across experiments with performance gains) achieved by NovelSeek and Dolphin. It can be observed that NovelSeek consistently improves the performance compared to the baseline and outperforms Dolphin across all tasks including both generative and discriminative tasks. This suggests that NovelSeek can generate better ideas on each specific domain benefiting from the self-evolving idea generation process and automatically implement them. For example, in AutoRYP, methods proposed by NovelSeek can largely outperform those proposed by Dolphin (i.e., +3.6 on max performance). We highlight that NovelSeek can achieve SoTA performance on some tasks such as 3D point cloud classification (i.e., 95.5% overall accuracy without pre-training achieved by NovelSeek compared to 95.3% overall accuracy with pre-training achieved by human experts).
@@
    Besides, Tab. 3 and Tab. 4 report the percentage of experiments with performance gains and executable experiments out of the total number of experiments. First, results show that even on complex tasks such as AutoPCDet (i.e., 50%) and Auto2DSeg (i.e., 90%), NovelSeek can still ensure a reasonable execution success rate which is due to the carefully designed idea-to-methodology process, enabling the coder to auto-implement based on detailed methodologies. Second, NovelSeek demonstrates a higher performance improvement rate compared to Dolphin. This improvement is mainly attributed to the idea-to-methodology feature of NovelSeek, which enables the concretization of high-level ideas. Additionally, through the process of multi-round experimental planning and execution, the submodules of the AI-generated methodology are progressively integrated into the baseline code.
@@
  • Covering a wide range of tasks including the scientific research tasks and AI tasks. Further, NovelSeek exhibits strong generalization capability across a wide range of tasks, enabling it to handle tasks from the AI domain (e.g., Auto2DSeg) to the scientific domain (e.g., AutoMD). As shown in Tab. 1 and Tab. 2, NovelSeek can support 12 different tasks ranging from simple classification tasks to complex multimodal and cross-disciplinary tasks. This is because the survey agent in NovelSeek can auto-search task-related literature on academic websites such as arXiv and review the literature to understand each task. Besides, NovelSeek is highly extensible, as it can support new tasks with just a task description and reference codes. This capability not only assists AI researchers in automatically updating algorithms, but also empowers researchers in scientific domains to utilize AI tools at a lower cost, thereby accelerating the pace of scientific discovery.
@@
  • Support repo-level experiments. Most of existing auto-research systems such as Dolphin (Yuan et al., 2025) only support single-file experiments. On more complex tasks, researchers are required to manually consolidate complex task codes into a single file, which is highly time-consuming and limits their ability to conduct experiments on complex tasks. In contrast, NovelSeek can support repo-level tasks such as AutoPCDet, AutoVLM, AutoTPPR, and so on, and achieve better performance on these repo-level tasks compared to their baselines. For example, on Auto2DSeg, NovelSeek pipeline can improve the DeepLabV3Plus baseline (Chen et al., 2018) from the original 78.80% to 81.0%. This is attributed to the detailed methodology, code comprehension achieved by the code review agent, and the auto-exploration ability of the coder agent.
@@
Runtime Statistics. We further provide the runtime statistics of NovelSeek on all 12 tasks including the training costs (i.e., GPU hours) and monetary costs in the idea generation stage (including self-evolving idea generation and idea-to-methodology) and code execution and debug stage. As shown in Tab. 5 and Tab. 6. As mentioned in Sec. 3.1.3, we select top 5 ideas in each idea generation process and then generate detailed methodology for the selected ideas. Therefore, we report the average cost of 5 ideas as the idea generation cost. It can be seen that the idea generation cost of each idea is about \$0.6 using GPT-4o which is cost-efficient. The coder-debug cost denote the cost of each run, for example, if running for 5 times for a single idea as mentioned in Sec. 2.3, we calculate the average cost of 5 runs. It can be seen from the table that the coder-debug cost varies between the file-level codes and repo-level codes and repo-level codes generally need more cost for high complexity of codes. For example, for single-file code such as Auto2DCls, the cost is below \$1 for each run and for more complex AutoPCDet, the cost is about \$1.2 using claude-sonnet-3.7. Generally, NovelSeek is a cost-efficient auto-research framework that can generate ideas and execute codes at a reasonable cost.
@@
Table 6: Computational and financial cost analysis for all tasks. Training time is measured using A100 GPU hours, while idea generation and code debugging costs are measured in USD using gpt-4o and claude-sonnet-3.7 models respectively.
@@
Coder-Debug cost (claude-sonnet-3.7) (\$) 0.7 0.6 0.7 1.1 1.2 1.0
Table 7: To compare the performance of the baseline and NovelSeek-generated code, we adopted a few-shot training setup for the yield prediction task. Due to the large variance in experimental results under this setting, we report the outcomes of 5 independent repeated experiments.

### v3

# InternAgent: When Agent Becomes the Scientist – Building Closed-Loop System from Hypothesis to Verification

@@ Abstract
Artificial Intelligence (AI) is accelerating the transformation of scientific research paradigms, not only enhancing research efficiency but also driving innovation. We introduce InternAgent, a unified closed-loop multi-agent framework to conduct Autonomous Scientific Research (ASR) across various scientific research fields, enabling researchers to tackle complicated problems in these fields with unprecedented speed and precision. InternAgent highlights three key advantages: 1) Scalability: InternAgent has demonstrated its versatility across 12 scientific research tasks, capable of generating innovative ideas to enhance the performance of baseline code. 2) Interactivity: InternAgent provides an interface for human expert feedback and multi-agent interaction in automated end-to-end processes, allowing for the seamless integration of domain expert knowledge. 3) Efficiency: InternAgent has achieved promising performance gains in several scientific fields with significantly less time cost compared to human efforts. For instance, in reaction yield prediction, it increased from 27.6% to 35.4% in just 12 hours; in enhancer activity prediction, the Pearson correlation coefficient rose from 0.65 to 0.79 with only 4 hours of processing; and in 2D semantic segmentation, precision advanced from 78.8% to 81.0% in a mere 30 hours.
Figure 1: InternAgent can support 12 types of scientific research tasks ranging from the AI field to the science field, including reaction yield prediction, molecular dynamics, power flow estimation, time series forecasting, transcription prediction, enhancer activity prediction, sentiment classification, 2D image classification, 3D point classification, 2D semantic segmentation, 3D autonomous driving, large vision-language model fine-tuning.
@@ 1 Introduction
To further facilitate the advancement of ASR, we propose the InternAgent, an end-to-end auto-research pipeline, which covers four main modules: self-evolving idea generation, human-interactive feedback, idea-to-methodology construction, and multi-round experiment planning and execution. With the help of the Self-Evolving Human-interactive Idea Generation and Idea-to-Methodology Construction, InternAgent can transform a rough proposal into a detailed and easily implementable method, which further increases the success rate of code implementation process and enhances the efficiency of closed-loop experiments. Besides, by leveraging multi-round experiment planning and execution, InternAgent is capable of designing experimental plans and decomposing the experimental process according to the InternAgent-proposed modules, thereby validating the effectiveness of each InternAgent-generated module through experimentation.
As shown in Fig. 1, InternAgent has been validated across 12 scientific research tasks, and we are excited to see that the experimental results demonstrate the significant value of InternAgent in the entire process from hypothesis generation to experimental validation. For instance, in the Reaction Yield Prediction task, the baseline model only achieved a performance of $24.2\%\pm 4.2$, while our model improved it to $34.8\%\pm 1.1$ in just 12 hours. In contrast, human researchers typically require several months to achieve a similar level of performance improvement. Another example of performance improvement is the Enhancer Activity Prediction task. The baseline model, DeepSTARR, achieved a result of 0.65. By utilizing InternAgent to search relevant domain literature, automatically generate code, and conduct validation, the performance can be improved to 0.79, representing a promising enhancement. In addition, InternAgent also supports complex project-level modifications and debugging, which consist of multiple code files. These results clearly indicate that InternAgent can autonomously generate ideas and design algorithms, effectively reducing the dependence on human effort in scientific research. To facilitate reproducibility, we have open-sourced both the baselines and the codes generated by InternAgent used in all involved scientific research tasks at https://github.com/Alpha-Innovator/InternAgent.
@@
  • Unified Multi-agent Framework for Diverse Scientific Research Tasks: We present InternAgent, a unified closed-loop scientific research framework that can automate the entire research cycle, including idea generation, idea-to-methodology transformation, experiment execution, and result feedback. This unified framework can be directly applied to various scientific research scenarios and fields.
  • Interactive Interfaces for Enhanced Cooperative Research: InternAgent offers interactive interfaces for human-machine collaboration within the idea generation module and across the entire system. By selecting collaboration modes, such as leveraging AI or human experts, it provides evaluations of idea generation effectiveness and facilitates the assessment, reflection, and documentation of experimental results.
  • Comprehensive Experimental Validation and Human Studies: We conducted extensive human studies centered around InternAgent, including inviting domain experts to evaluate and score the novelty of ideas generated by InternAgent, and comparing the research efficiency between human researchers and InternAgent. These experiments and human studies are crucial for gaining insights into the capabilities of multi-agent systems in conducting scientific research tasks in open-ended environments. We observed many promising phenomena, while also identifying certain technical modules that require improvement.
Figure 2: InternAgent covers three main capabilities: 1) Self-evolving Idea Generation with Human-interactive Feedback, 2) Idea-to-Methodology Construction, and 3) Evolutionary Experimental Planning and Execution.
2 InternAgent
As a unified closed-loop multi-agent framework for ASR, InternAgent is designed to facilitate innovative research across diverse scientific domains, as illustrated in Fig. 2. It incorporates three primary capabilities: self-evolving idea generation with human-interactive feedback (Sec. 2.1), comprehensive idea-to-methodology construction (Sec. 2.2), and multi-round automated experiment execution (Sec. 2.3). Each capability is realized through the collaboration of specialized agents, allowing for seamless integration of different processes to enhance scientific discoveries.
@@ 2.1 Self-Evolving Idea Generation with Human-interactive Feedback
The self-evolving idea generation capability is central to InternAgent, enabling the framework to autonomously generate and refine innovative research ideas. This process involves several specialized agents, each contributing to different stages of idea development and refinement.
@@
By dynamically adjusting its search strategies based on the context of the research stage, the Survey Agent ensures a comprehensive and nuanced understanding of the research landscape. This capability not only supports the generation of innovative ideas but also ensures that the InternAgent framework remains at the cutting edge of scientific discovery.
Code Review Agent. The Code Review Agent is crucial for understanding baseline codes for different research tasks, serving as a foundation for innovation by identifying improvements and developing new methodologies. It provides detailed analyses of code structures, dependencies, and functionalities, enabling InternAgent to fully comprehend existing code-bases and identify potential enhancements to advance research objectives. Moreover, the agent’s ability to document and summarize complex code-bases ensures efficient navigation and utilization of existing methods. The agent manages two scenarios: 1) reviewing user-provided code or 2) searching for relevant code-bases. For user-uploaded code, it conducts a comprehensive review of the structure, logic, and functionality. Alternatively, in the absence of user-uploaded code, it searches public repositories like GitHub to find relevant code-bases, performing thorough analyses at both the repository and file levels to understand inter-dependencies and assess logic, efficiency, and correctness. Furthermore, the agent employs static code analysis using Python’s ‘ast‘ module to parse and understand code structure without execution, while the LLM generates human-readable descriptions and summaries, transforming technical details into structured documentation. By using parallel processing with Python’s ‘multiprocessing‘ module, the agent enhances efficiency and scalability for large code-bases. Overall, the Code Review Agent offers detailed documentation that deepens the understanding of code repositories and supports innovation in scientific research.
Idea Innovation Agent. The Idea Innovation Agent is an integral part of InternAgent, designed to enhance the creative and iterative processes of scientific research. This agent plays a crucial role by automating the generation and evolution of ideas, thereby addressing the limitations of traditional research works (Yuan et al., 2025; Yamada et al., 2025; Lu et al., 2024), which often rely on time-consuming manual efforts and are constrained by human cognitive biases. The agent’s dual responsibilities, idea generation and idea evolution, are specifically designed to address the diverse needs of various scientific disciplines.
@@
Overall, the Idea Innovation Agent enhances scientific ideas into viable and creative solutions by synthesizing and contextualizing information. It critically examines current ideas and employs feedback loops with human experts and other InternAgent agents for continuous improvement. This iterative process balances novelty, feasibility, and ethical considerations, producing impactful and well-rounded ideas.
Assessment Agent. The Assessment Agent is a vital component of InternAgent, designed to ensure the quality and viability of generated ideas through a rigorous evaluation process. In the rapidly evolving landscape of scientific research, the systematic and objective assessment of ideas is essential. Traditional methods often suffer from subjectivity and lack comprehensive coverage of all relevant dimensions, which can lead to promising ideas being overlooked(Qiu et al., 2025; Si et al., 2024). Therefore, the Assessment Agent addresses these challenges by providing a structured and multidimensional evaluation process, which in turn enhances the reliability and effectiveness of idea selection.
@@
The human-interactive feedback mechanism of InternAgent is categorized into two primary types: 1) feedback directly provided by humans and 2) feedback automatically generated by agent. Human-provided feedback can address one or multiple ideas, offering insights and critiques that lead to further refinement and adjustment of these ideas based on the feedback received. This iterative process facilitates the continuous improvement of ideas, ensuring they are honed to meet specific objectives and challenges.
@@
In summary, as illustrated in Fig. 3, by managing multi-agent collaboration and integrating human feedback, the Orchestration Agent enables InternAgent to operate as a cohesive and innovative research tool, driving scientific discovery forward.
Figure 3: InternAgent Self-evolutionary path of ideas for reaction yield prediction task.
@@ 2.2.2 Methodology Refinement
During both initialization and refinement, the Methodology Development Agent collaborates closely with other agents, such as the Assessment Agent for multidimensional evaluation and the Orchestration Agent for workflow coordination. This collaboration ensures that each methodological step benefits from comprehensive feedback and current domain knowledge. The integrated, multi-agent approach guarantees that the transformation from idea to methodology is systematic and adaptable, supporting the continuous evolution and optimization of scientific research within the InternAgent framework.
@@ 3 Experiments
In this section, we evaluate the effectiveness of InternAgent in conducting autonomous research and accelerating scientific discovery. We begin by providing a brief overview of the selected multi-domain tasks and detailing the experimental implementation in Sec. 3.1. Subsequently, we present the quantitative results across various tasks in Sec. 3.2 and conduct an analysis of the different modules within InternAgent in Sec. 3.3.
@@ 3.1.1 Task Description
We select 12 distinct tasks to demonstrate InternAgent’s capability in conduct Autonomous Scientific Research (ASR). These tasks span multiple modalities, including science (e.g., reaction yield prediction, molecular dynamics), time series (e.g., time series forecasting), natural language (e.g., sentiment classification), image (e.g., semantic segmentation), and point cloud (e.g., 3D object detection), which cover both discriminative and generative tasks. We believe that experiments ranging from fundamental tasks to complex multi-modal tasks can comprehensively illustrate the effectiveness of InternAgent. Below, we detail the datasets, the base code repositories, and the experimental settings for each task.
@@ 3.1.2 Evaluation Metric
Since our InternAgent has been validated across a wide range of scientific research fields, the evaluation metrics used for tasks in each field are not consistent. Therefore, in this part, we provide a detailed introduction to the evaluation metrics used for each scientific research task.
@@ 3.2 Experimental Results
Table 1: Performance comparison across six types of scientific research tasks. We conduct experiments using 10 InternAgent-generated ideas for each task.
@@
InternAgent 33.5 (+5.9) 0.152 0.00447 0.4346 0.170 0.77
Table 2: Performance comparison for six types of scientific research tasks. We conduct experiments using 10 InternAgent-generated ideas for each task, where baseline codes for Auto2DSeg, AutoPCDet, and AutoVLM are project-level, consisting of multiple code files with complex call relation between functions. Therefore, the coder in Dolphin (Yuan et al., 2025) does not support modifying this type of baseline codes.
@@
InternAgent 5/7/10 3/6/10 9/9/10 6/9/10 2/5/10 1/5/10
To comprehensively evaluate the effectiveness of InternAgent in accelerating scientific discovery, we first provide quantitative experimental results as shown in Tab. 1, Tab. 2, Tab. 3, and Tab. 4. Extensive results demonstrate that InternAgent excels in the following aspects:
Table 5: Computational and financial cost analysis for all tasks. Training time is measured using A100 GPU hours, while idea generation and code debugging costs are measured in USD using gpt-4o and claude-3.7-sonnet models respectively.
@@
Coder-Debug cost (claude-3.7-sonnet) (\$) 0.7 0.5 1.0 0.4 0.9 0.6
@@
  • Outperforming existing auto-research systems on multiple tasks. We first compare InternAgent with existing auto-research system (i.e., Dolphin (Yuan et al., 2025)) on single-file tasks. Tab. 1 and Tab. 2 show the max performance and average performance (i.e., the average performance across experiments with performance gains) achieved by InternAgent and Dolphin. It can be observed that InternAgent consistently improves the performance compared to the baseline and outperforms Dolphin across all tasks including both generative and discriminative tasks. This suggests that InternAgent can generate better ideas on each specific domain benefiting from the self-evolving idea generation process and automatically implement them. For example, in AutoRYP, methods proposed by InternAgent can largely outperform those proposed by Dolphin (i.e., +3.6 on max performance). We highlight that InternAgent can achieve SoTA performance on some tasks such as 3D point cloud classification (i.e., 95.5% overall accuracy without pre-training achieved by InternAgent compared to 95.3% overall accuracy with pre-training achieved by human experts).
@@
    Besides, Tab. 3 and Tab. 4 report the percentage of experiments with performance gains and executable experiments out of the total number of experiments. First, results show that even on complex tasks such as AutoPCDet (i.e., 50%) and Auto2DSeg (i.e., 90%), InternAgent can still ensure a reasonable execution success rate which is due to the carefully designed idea-to-methodology process, enabling the coder to auto-implement based on detailed methodologies. Second, InternAgent demonstrates a higher performance improvement rate compared to Dolphin. This improvement is mainly attributed to the idea-to-methodology feature of InternAgent, which enables the concretization of high-level ideas. Additionally, through the process of multi-round experimental planning and execution, the submodules of the AI-generated methodology are progressively integrated into the baseline code.
@@
  • Covering a wide range of tasks including the scientific research tasks and AI tasks. Further, InternAgent exhibits strong generalization capability across a wide range of tasks, enabling it to handle tasks from the AI domain (e.g., Auto2DSeg) to the scientific domain (e.g., AutoMD). As shown in Tab. 1 and Tab. 2, InternAgent can support 12 different tasks ranging from simple classification tasks to complex multimodal and cross-disciplinary tasks. This is because the survey agent in InternAgent can auto-search task-related literature on academic websites such as arXiv and review the literature to understand each task. Besides, InternAgent is highly extensible, as it can support new tasks with just a task description and reference codes. This capability not only assists AI researchers in automatically updating algorithms, but also empowers researchers in scientific domains to utilize AI tools at a lower cost, thereby accelerating the pace of scientific discovery.
@@
  • Support repo-level experiments. Most of existing auto-research systems such as Dolphin (Yuan et al., 2025) only support single-file experiments. On more complex tasks, researchers are required to manually consolidate complex task codes into a single file, which is highly time-consuming and limits their ability to conduct experiments on complex tasks. In contrast, InternAgent can support repo-level tasks such as AutoPCDet, AutoVLM, AutoTPPR, and so on, and achieve better performance on these repo-level tasks compared to their baselines. For example, on Auto2DSeg, InternAgent pipeline can improve the DeepLabV3Plus baseline (Chen et al., 2018) from the original 78.80% to 81.0%. This is attributed to the detailed methodology, code comprehension achieved by the code review agent, and the auto-exploration ability of the coder agent.
@@
Runtime Statistics. We further provide the runtime statistics of InternAgent on all 12 tasks including the training costs (i.e., GPU hours) and monetary costs in the idea generation stage (including self-evolving idea generation and idea-to-methodology) and code execution and debug stage. As shown in Tab. 5 and Tab. 6. As mentioned in Sec. 3.1.3, we select top 5 ideas in each idea generation process and then generate detailed methodology for the selected ideas. Therefore, we report the average cost of 5 ideas as the idea generation cost. It can be seen that the idea generation cost of each idea is about \$0.6 using GPT-4o which is cost-efficient. The coder-debug cost denote the cost of each run, for example, if running for 5 times for a single idea as mentioned in Sec. 2.3, we calculate the average cost of 5 runs. It can be seen from the table that the coder-debug cost varies between the file-level codes and repo-level codes and repo-level codes generally need more cost for high complexity of codes. For example, for single-file code such as Auto2DCls, the cost is below \$1 for each run and for more complex AutoPCDet, the cost is about \$1.2 using claude-3.7-sonnet. Generally, InternAgent is a cost-efficient auto-research framework that can generate ideas and execute codes at a reasonable cost.
@@
Table 6: Computational and financial cost analysis for all tasks. Training time is measured using A100 GPU hours, while idea generation and code debugging costs are measured in USD using gpt-4o and claude-3.7-sonnet models respectively.
@@
Coder-Debug cost (claude-3.7-sonnet) (\$) 0.7 0.6 0.7 1.1 1.2 1.0
Table 7: To compare the performance of the baseline and InternAgent-generated code, we adopted a few-shot training setup for the yield prediction task. Due to the large variance in experimental results under this setting, we report the outcomes of 5 independent repeated experiments.